OpenStudy (anonymous):

Use the explicit formula an = a1 + (n - 1) * d to find the 1000th term of the sequence below. 24, 31, 38, 45, 52, ... A. 7024 B. 7017 C. 6993 D. 7045 Would it be B?

OpenStudy (anonymous):

a1=21 d=a2-a1=? n=1000 calculate an

OpenStudy (anonymous):

ok im kinda lost?

OpenStudy (freckles):

first what is your common difference @tallan ?

OpenStudy (freckles):

@surjithayer was saying to find the common difference take a term from your sequence and subtract the previous term can you do that?

OpenStudy (freckles):

for example what is 31-24 or 38-31 or 45-38 or 52-45 these differences should all be the same since this is arithmetic sequence so you only need to find one of those differences

OpenStudy (freckles):

to determine the value of d

OpenStudy (anonymous):

i got it thanks!! The answer is B.. But I do have another question... could you help me on it?

OpenStudy (freckles):

sure I can try

OpenStudy (anonymous):

Which of the following choices is the simple formula for the nth term of the following arithmetic sequence? 5, 1, -3, -7, -11, ...

OpenStudy (anonymous):

A. 4n + 1 B. 4n + 9 C. 5n + 4 D. -4n + 1 E. -4n + 9

OpenStudy (freckles):

well let's look at the form \[a_n=a_1+d(n-1)\] d is the common difference a_1 is the first term of your sequence -- to find d; just take a term and subtract its previous term a_1 should be pretty easy to see (After all it is the first number in the sequence)

OpenStudy (freckles):

we will enter these two numbers into the formula above

OpenStudy (freckles):

then do a little distributing and combining like terms

OpenStudy (freckles):

\[a_n=a_1+dn-d \\ a_n=dn+(a_1-d)\] that is you will be able to find a1-d

OpenStudy (freckles):

have you found the common difference yet?

OpenStudy (anonymous):

im really slow at math so im trying to figure it out @freckles

OpenStudy (freckles):

again to find the common difference take a term and then subtract it's previous term so you perform anyone of the differences: 1-5 -3-1 -7-(-3) -11-(-7)

OpenStudy (anonymous):

-4?

OpenStudy (freckles):

so replace d with -4 and replace a1 with 5

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

9?

OpenStudy (freckles):

\[a_n=a_1+dn-d \\ a_n=dn+(a_1-d) \\ a_n=-4n+9\] a1-d is definitely 9 :)

OpenStudy (anonymous):

Thank you so much for your help!!!

OpenStudy (freckles):

np